# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v2

@@ Abstract
Graph generation is a critical yet challenging task as empirical analyses require a deep understanding of complex, non-Euclidean structures. Diffusion models have recently made significant achievements in graph generation, but these models are typically adapted from image generation frameworks and overlook inherent higher-order topology, leaving them ill-suited for capturing the topological properties of graphs. In this work, we propose Higher-order Guided Diffusion (HOG-Diff), a principled framework that progressively generates plausible graphs with inherent topological structures. HOG-Diff follows a coarse-to-fine generation curriculum guided by higher-order topology and implemented via diffusion bridges. We further prove that our model exhibits a stronger theoretical guarantee than classical diffusion frameworks. Extensive experiments on both molecular and generic graph generation tasks demonstrate that our method consistently outperforms or remains competitive with state-of-the-art baselines. Our code is available at https://github.com/Yiminghh/HOG-Diff.
@@ 1 Introduction
Figure 1: Overview of HOG-Diff. The dashed line above illustrates the classical generation process, where graphs quickly degrade into random structures with uniformly distributed entries. In contrast, as shown in the coloured region below, HOG-Diff adopts a coarse-to-fine generation curriculum based on the diffusion bridge, explicitly learning higher-order structures during intermediate steps with theoretical guarantees.
Graphs provide an elegant abstraction for representing complex systems by encoding entities as vertices and their relationships as pairwise edges. Yet mounting evidence shows that many empirical systems are governed by higher-order group interactions (Battiston et al., 2020). For instance, many molecules must assemble into specific structures to function properly, academic collaborations are most productive at specific group scales, and neural activity emerges from coordinated population dynamics (Papamarkou et al., 2024). Recent studies further reveal that higher-order interactions are ubiquitous and essential for understanding complex phenomena in domains as diverse as online social dynamics, ecological stability, and neuronal systems (Gardner et al., 2022).
Graph generation seeks to synthesize novel graphs faithful to the observed data distribution, with applications spanning drug discovery, motion synthesis, social network analysis, and urban planning (Zhu et al., 2022). However, modeling the underlying distribution of relational data remains challenging, partly because conventional approaches often overlook or implicitly learn data topology. Since higher-order structures are intrinsic to many empirical systems, explicitly incorporating them into generative frameworks promises to simplify generation and enhance realism. For instance, statistical analyses reveal that approved drug molecules occupy only a few hundred distinct higher-order structures (Ertl et al., 2025), in sharp contrast to the astronomically large chemical space of drug-like molecules ($10^{23}$$10^{60}$(Segler et al., 2018), suggesting that leveraging such topological guides could yield more efficient molecular design. Recent advances in topological deep learning (TDL) further reinforce this view, demonstrating that higher-order topology can substantially improve graph and molecular representation learning (Hajij et al., 2022; Papamarkou et al., 2024; Liu et al., 2024; Battiloro et al., 2025). Yet, despite its promise, higher-order topology has not been explicitly integrated into generative models, leaving a significant gap.
Early work on graph generative models dates back to random network models (Barabási & Albert, 1999), which offer foundational insights but are too simplistic for capturing real-world graph distributions. Recent advances in generative models have leveraged the power of deep neural networks, significantly improving the ability to learn graph distributions. Notable approaches include models based on recurrent neural networks (RNNs) (You et al., 2018), variational autoencoders (VAEs) (Jin et al., 2018), and generative adversarial networks (GANs) (Martinkus et al., 2022). However, the end-to-end structure of these methods makes them hard to train. More recently, diffusion-based models have achieved remarkable success in image generation by learning a model to denoise a noisy sample (Ho et al., 2020; Song et al., 2021; Yue et al., 2024). With the advent of diffusion models, their applications on graphs with complex topological structural properties have recently aroused significant scientific interest (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2023; Qin et al., 2025). Despite these advances, diffusion-based generative approaches are ineffective at modeling the topological properties of higher-order systems since learning to denoise the noisy samples does not explicitly preserve the intricate structural dependencies required for generating realistic graphs.
Moreover, existing diffusion-based models for graphs typically inherit the frameworks designed for image generation (Song et al., 2021), which fundamentally limits their ability to capture the intrinsic topological properties of networks. Specifically, the image corrupted by Gaussian noise retains recognizable patterns during the early and middle stages of forward diffusion. By contrast, the graph adjacency matrix quickly degrades into a dense matrix with uniformly distributed entries within a few diffusion steps. In addition, applying diffusion-based generative models directly to graph topology generation by injecting isotropic Gaussian noise into adjacency matrices is detrimental, as it destroys essential graph properties such as sparsity and connectivity (Yang et al., 2023). Lastly, such a framework should ensure equivariance, with invariance as a special case, so that the learned distribution remains unchanged under node reordering, which is essential for robustness and capturing intrinsic graph distribution. Taken together, these challenges highlight the need for a graph-friendly diffusion framework that explicitly learns higher-order topology, preserves meaningful intermediate states and trajectories, avoids inappropriate noise addition, and ensures equivariance.
Motivated by these principles, we propose the Higher-order Guided Diffusion (HOG-Diff) framework (see Fig. 1), a coarse-to-fine generation curriculum that preserves higher-order topologies throughout the diffusion process to better capture complex graph structures. Specifically, we decompose the graph generation task into manageable sub-tasks, beginning by generating higher-order skeletons that capture core structures, which are then refined to include pairwise interactions and finer details, ultimately synthesizing complete graphs with both topological and semantic fidelity. Additionally, HOG-Diff integrates the generalized Ornstein-Uhlenbeck (OU) bridge in the spectral domain to ensure effectiveness and adherence to the aforementioned graph generation principles. Our theoretical analysis reveals that HOG-Diff converges more rapidly in score matching and achieves sharper reconstruction error bounds than classical approaches, offering strong theoretical support for the proposed framework. Furthermore, our framework promises to enhance interpretability by enabling the analysis of different topological guides’ performance in the generation process. The contributions of this paper are fourfold:
@@
  • We introduce cell complex filtering to extract higher-order skeletons as valuable generation guides.
@@
  • We propose a principled coarse-to-fine graph generation framework guided by higher-order topological information and implemented via the generalized OU diffusion bridge.
@@
  • Extensive evaluations on molecules and generic graph datasets show that HOG-Diff achieves state-of-the-art performance, highlighting the practical importance of topological guidance.
@@ 2 Preliminaries
Graphs are elegant and useful abstractions for various empirical objects. Formally, a graph can be represented as $\bm{G}\triangleq(\bm{V},\bm{E},\bm{X})$, where $\bm{V}$ denotes the node set, $\bm{E}\subseteq\bm{V}\times\bm{V}$ the edges, and $\bm{X}$ the nodes feature matrix. However, many empirical systems exhibit group interactions that go beyond simple pairwise relationships (Battiston et al., 2020). To capture these complex interactions, higher-order networks—such as hypergraphs, simplicial complexes (SCs), and cell complexes (CCs)—offer more expressive alternatives by capturing higher-order interactions among multiple entities (Papamarkou et al., 2024). Among these, cell complexes are fundamental in algebraic topology, offering a flexible generalization of pairwise graphs (Hatcher, 2001).
@@
A fundamental goal of generative models is to produce plausible samples from an unknown target data distribution $p(\mathbf{x}_{0})$. Score-based diffusion models (Song & Ermon, 2019; Song et al., 2021) achieve this by progressively corrupting the authentic data with noise and subsequently training a neural network to reverse this corruption process, thereby generating meaningful data from a tractable prior distribution, i.e., $\mathbf{x}_{\mathrm{generated}}\sim p(\mathbf{x}_{0})$.
@@
Starting from time $T$, $p(\mathbf{x}_{T})$ can be progressively transformed back to $p(\mathbf{x}_{0})$ by following the trajectory of the reverse-time SDE $\mathrm{d}\mathbf{x}_{t}=[\mathbf{f}_{t}(\mathbf{x}_{t})-g_{t}^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})]\mathop{}\!\mathrm{d}{\bar{t}}+g_{t}\mathrm{d}\bar{\mathbf{w}}_{t}$ (Anderson, 1982), where $p_{t}(\cdot)$ denotes the probability density function of $\mathbf{x}_{t}$ and $\bar{\mathbf{w}}$ is a reverse-time Wiener process.
The term $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, known as the score function, is typically parameterized by a neural network $\bm{s}_{\bm{\theta}}(\mathbf{x}_{t},t)$ and trained using the conditional score-matching loss function (Vincent, 2011):
@@
Doob’s $h$-transform is a mathematical framework widely used to modify stochastic processes, enabling the process to satisfy specific terminal conditions. By introducing a $h$-function into the drift term of an SDE, this technique ensures that the process transitions to a predefined endpoint while preserving the underlying probabilistic structure. Specifically, given the SDE in Eq. 1, Doob’s $h$-transform alters the SDE to include an additional drift term, ensuring that the process reaches a fixed terminal state at $t=T$. The modified SDE is expressed as:
@@ 3.1 Coarse-to-fine Framework with Topological Filtering
To implement our coarse-to-fine generation framework, we first introduce a key operation termed cell complex filtering (CCF). The filtering operation decomposes the graph generation task into hierarchically structured and manageable subtasks.
@@
Given a graph $\bm{G}=(\bm{V},\bm{E})$ and its associated cell complex $\mathcal{S}=\cup_{\alpha}x_{\alpha}$ (obtained via lifting). The $p$-cell complex filtering operation defines a filtered graph $\bm{G}_{p}=(\bm{V}_{p},\bm{E}_{p})$, where $\bm{V}_{p}=\{v\in\bm{V}\mid\exists\;x_{\alpha}\text{with}\dim(x_{\alpha})=p:v\in\overline{x_{\alpha}}\}$, and $\bm{E}_{p}=\{(u,v)\in\bm{E}\mid\exists\;x_{\alpha}\text{with}\dim(x_{\alpha})=p:\{u,v\}\subseteq\overline{x_{\alpha}}\}$. Here, $\overline{x_{\alpha}}$ denotes the closure of $x_{\alpha}$.
As illustrated in Fig. 2, we first lift the graph to a cell complex and then apply CCF to generate intermediate states by pruning nodes and edges that do not belong to higher-order cells. In practice, CCF offers a substantial speedup, as it avoids the expensive enumeration of all cells required by lifting, see App. D for details. Based on the filtering results, the diffusion process is structured into $K$ hierarchical time windows, denoted as $\{[\tau_{k-1},\tau_{k}]\}_{k=1}^{K}$, where $0=\tau_{0}<\cdots<\tau_{k-1}<\tau_{k}<\cdots<\tau_{K}=T$, with the filtered results serving as natural intermediaries in the hierarchical generation process.
The overall framework of HOG-Diff is depicted in Fig. 1. In general, we first generate coarse-grained higher-order skeletons and subsequently refine them into finer pairwise relationships, thereby simplifying the task of capturing complex graph distributions. Formally, our generation process factorizes the joint distribution of the final graph $\bm{G}_{0}$ into a product of conditional distributions across these time windows:
$p(\bm{G}_{0})=p(\bm{G}_{0}|\bm{G}_{\tau_{1}})p(\bm{G}_{\tau_{1}}|\bm{G}_{\tau_{2}})\cdots p(\bm{G}_{\tau_{K-1}}|\bm{G}_{T}).$
@@
$\mathrm{d}\bm{G}_{t}^{(k)}=\mathbf{f}_{k,t}(\bm{G}_{t}^{(k)})\mathrm{d}t+g_{k,t}\mathrm{d}\bm{W}_{t},t\in[\tau_{k-1},\tau_{k}].$ (3)
@@ 3.2 Guided Generation via Diffusion Bridge Process
Applying Doob’s $h$-transform (Doob & Doob, 1984) to the GOU process under the terminal condition $\bm{\mu}=\bm{G}_{\tau_{k}}$, we can derive the GOU bridge process as follows (we provide the detailed derivation of the bridge process in Sec. A.1):
@@
where $\omega(t)$ is a positive weighting function, and $c_{1},c_{2}$ controls the relative importance of vertices and spectrum. The training procedure is detailed in Alg. 1 in App. C.
@@
To achieve this, we introduce a unified score network that explicitly integrates node and spectral representations. As illustrated in Fig. 7 of Appendix, the network comprises two different graph processing modules: a standard graph convolution network (GCN) (Kipf & Welling, 2017) for local feature aggregation and a graph transformer network (ATTN) (Dwivedi & Bresson, 2021; Vignac et al., 2023) for global information extraction. The outputs of these modules are fused with time information through a Feature-wise Linear Modulation (FiLM) layer (Perez et al., 2018), and the resulting representations are concatenated to form a unified hidden embedding. This hidden embedding is further processed by separate multilayer perceptrons (MLPs) to produce predictions for $\nabla_{\bm{X}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$ and $\nabla_{\bm{\Lambda}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$, respectively. It is worth noting that our score network is permutation equivalent, as each component of our model avoids any node ordering-dependent operations. Our model is detailed in Sec. C.3.

### v3

@@ Abstract
Graph generation is a critical yet challenging task, as empirical analyses require a deep understanding of complex, non-Euclidean structures. Diffusion models have recently made significant advances in graph generation, but these models are typically adapted from image generation frameworks and overlook inherent higher-order topology, limiting their ability to capture graph topology. In this work, we propose Higher-order Guided Diffusion (HOG-Diff), a principled framework that progressively generates plausible graphs with inherent topological structures. HOG-Diff follows a coarse-to-fine generation curriculum, guided by higher-order topology and implemented via diffusion bridges. We further prove that our model admits stronger theoretical guarantees than classical diffusion frameworks. Extensive experiments across eight graph generation benchmarks, spanning diverse domains and including large-scale settings, demonstrate the scalability of our method and its superior performance on both pairwise and higher-order topological metrics. Our project page is available here.
@@ 1 Introduction
Graphs provide an elegant abstraction for representing complex systems by encoding entities as vertices and their relationships as pairwise edges. As such, they have played a key role in generative modeling in unstructured domains, enabling the synthesis of novel graphs faithful to the data distribution. This representational power has positioned graph generative models as a crucial tool for discovering new molecules, materials, and biostructures (Jumper et al., 2021).
Despite these advances, graph generation still lags behind its Euclidean counterparts. Classically, oversmoothing and oversquashing limit expressive capacity (Alon and Yahav, 2021), but an even deeper obstacle persists: most generative frameworks treat graphs purely as collections of pairwise edges, overlooking the higher-order structures that govern the organization of real-world systems, such as triangles, cliques, rings, and motifs (Battiston et al., 2020). These higher-order structures play a decisive role in fields ranging from chemistry to neuroscience: molecules function through coordinated multi-atom assemblies; collaborative teams operate effectively only at specific group sizes; and neural populations exhibit coherent activity patterns that fundamentally rely on multi-unit interactions (Gardner et al., 2022; Papamarkou et al., 2024). A growing body of empirical and theoretical work underscores that such higher-order topology is not an optional structure, but a defining characteristic of real data (Segler et al., 2018; Ertl et al., 2025).
Recent studies make this picture even sharper. Approved drug molecules, for instance, inhabit only a few hundred distinct ring systems (Ertl et al., 2025), far fewer than the astronomical chemical space of plausible compounds ($10^{23}$$10^{60}$(Segler et al., 2018). Meanwhile, advances in topological deep learning (TDL) show that explicitly modeling complex structures boosts the expressivity and stability of graph and molecular representation learning (Hajij et al., 2022; Papamarkou et al., 2024; Liu et al., 2024; Wang et al., 2025; Hajij et al., 2025). Nevertheless, no existing graph generative model integrates higher-order structure as an explicit guiding signal.
@@
Figure 1: Overview of the HOG-Diff framework and its performance. (Left) The dashed trajectory above illustrates the classical generation process, where graphs rapidly degrade into structureless randomness. In contrast, as shown in the coloured region below, HOG-Diff adopts a coarse-to-fine generation curriculum via a diffusion bridge, explicitly learning higher-order structures during intermediate steps. (Middle and right) Relative improvements over prior state-of-the-art on pairwise and higher-order topological metrics.
In this work, motivated by this gap, we propose the Higher-order Guided Diffusion (HOG-Diff) framework, a principled, topology-aware framework that places higher-order topology at the core of the generative process. As illustrated in Fig. 1, HOG-Diff implements a coarse-to-fine generation curriculum: it first synthesizes the higher-order skeleton of a graph, its key 2-cells, faces, or other motifs, and then refines these coarse structures into full pairwise connectivity. To realize this, we combine cell-complex filtering with a novel generalized Ornstein–Uhlenbeck (OU) diffusion bridge in the spectral domain, enabling the diffusion process to transition smoothly between increasingly detailed topological states while preserving the global structure implied by higher-order interactions.
This coordinated two-stage design brings several benefits. It aligns the generative trajectory with the intrinsic hierarchical organization of real-world graphs; it avoids the collapse of intermediate states into meaningless noisy adjacency matrices; and it allows us to leverage closed-form bridge dynamics for stable, simulation-free training. Theoretically, we prove that HOG-Diff enjoys faster convergence in score matching and tighter reconstruction error bounds than classical diffusion models. Practically, the explicit presence of topological guides enables interpretability: by varying the guide structures, we can directly probe which topological motifs are most influential in determining the generative process. Across molecular and generic graph benchmarks, HOG-Diff consistently achieves state-of-the-art performance, sharply reducing both statistical and topological discrepancies relative to real data. Our findings underscore a central message: higher-order topology is a powerful generative signal and incorporating it transforms graph diffusion from an edge-level denoising procedure into a truly structure-aware generative paradigm. Our concrete contributions are:
@@
  • We introduce cell complex filtering to extract higher-order skeletons from graphs as valuable generation guides.
@@
  • We propose a principled, coarse-to-fine graph generation framework guided by higher-order topology and implemented via the generalized OU diffusion bridge.
@@
  • Comprehensive evaluations across eight benchmarks show that topology-informed HOG-Diff achieves state-of-the-art performance on both pairwise and higher-order metrics, highlighting the value of topological guidance.
@@ 2 Preliminaries
Graphs are elegant and useful abstractions for various empirical objects. Formally, a graph can be represented as $\bm{G}\triangleq(\bm{V},\bm{E},\bm{X})$, where $\bm{V}$ denotes the node set, $\bm{E}\subseteq\bm{V}\times\bm{V}$ the edges, and $\bm{X}$ the node feature matrix. However, many empirical systems exhibit group interactions that go beyond simple pairwise relationships (Battiston et al., 2020). To capture these complex interactions, higher-order networks—such as hypergraphs, simplicial complexes (SCs), and cell complexes (CCs)—offer more expressive alternatives by capturing higher-order interactions among multiple entities (Papamarkou et al., 2024). Among these, cell complexes are fundamental in algebraic topology, offering a flexible generalization of pairwise graphs (Hatcher, 2001).
@@
A fundamental goal of generative models is to produce plausible samples from an unknown target data distribution $p(\mathbf{x}_{0})$. Score-based diffusion models (Song and Ermon, 2019; Song et al., 2021) achieve this by progressively corrupting the authentic data with noise and subsequently training a neural network to reverse this corruption process, thereby generating meaningful data from a tractable prior distribution, i.e., $\mathbf{x}_{\mathrm{generated}}\sim p(\mathbf{x}_{0})$.
@@
Starting from time $T$, $p(\mathbf{x}_{T})$ can be progressively transformed back to $p(\mathbf{x}_{0})$ by following the trajectory of the reverse-time SDE $\mathrm{d}\mathbf{x}_{t}=[\mathbf{f}_{t}(\mathbf{x}_{t})-g_{t}^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})]\mathop{}\!\mathrm{d}{\bar{t}}+g_{t}\mathrm{d}\bar{\mathbf{w}}_{t}$ (Anderson, 1982), where $p_{t}(\cdot)$ denotes the probability density function of $\mathbf{x}_{t}$ and $\bar{\mathbf{w}}$ is a reverse-time Wiener process. The score function $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ is typically parameterized by a neural network $\bm{s}_{\bm{\theta}}(\mathbf{x}_{t},t)$ and trained using the conditional score-matching loss function (Vincent, 2011):
@@
Doob’s $h$-transform is a mathematical framework widely used to modify stochastic processes, enabling the process to satisfy specific terminal conditions. By introducing an $h$-function into the drift term of an SDE, this technique ensures that the process transitions to a predefined endpoint while preserving the underlying probabilistic structure. Specifically, given the SDE in Eq. 1, Doob’s $h$-transform alters the SDE to include an additional drift term, ensuring that the process reaches a fixed terminal state at $t=T$. The modified SDE is expressed as:
@@ 3.1 Coarse-to-fine Framework with Topological Filtering
To implement our coarse-to-fine generation, we first introduce a key operation termed cell complex filtering (CCF).
@@
Given a graph $\bm{G}=(\bm{V},\bm{E})$ and its associated cell complex $\mathcal{S}=\cup_{\alpha}x_{\alpha}$ (obtained via lifting). The $p$-cell complex filtering operation defines a filtered graph $\bm{G}_{[p]}=(\bm{V}_{[p]},\bm{E}_{[p]})$, where $\bm{V}_{[p]}=\{v\in\bm{V}\mid\exists\;x_{\alpha}\text{with}\dim(x_{\alpha})=p:v\in\overline{x_{\alpha}}\}$, and $\bm{E}_{[p]}=\{(u,v)\in\bm{E}\mid\exists\;x_{\alpha}\text{with}\dim(x_{\alpha})=p:\{u,v\}\subseteq\overline{x_{\alpha}}\}$. Here, $\overline{x_{\alpha}}$ denotes the closure of $x_{\alpha}$.
As illustrated in Fig. 2, we first lift the graph to a cell complex and then apply CCF to generate intermediate states by pruning nodes and edges that do not belong to higher-order cells. In practice, CCF offers a substantial speedup, as it avoids the expensive enumeration of all cells required by lifting (see App. D for detailed complexity analysis). The filtering operation decomposes the graph generation task into hierarchically structured and manageable subtasks.
Based on the filtering results, the diffusion process is structured into $K$ hierarchical time windows, denoted as $\{[\tau_{k-1},\tau_{k}]\}_{k=1}^{K}$, where $0=\tau_{0}<\cdots<\tau_{k-1}<\tau_{k}<\cdots<\tau_{K}=T$, with the filtered results serving as natural intermediaries in the hierarchical generation process. The overall framework of HOG-Diff is depicted in Fig. 1(left). In general, we first generate coarse-grained higher-order skeletons and subsequently refine them into finer pairwise relationships, thereby simplifying the task of capturing complex graph distributions. Formally, our generation process factorizes the joint distribution of the final graph $\bm{G}_{0}$ into a product of conditional distributions across these time windows:
$p(\bm{G}_{0})=p(\bm{G}_{0}|\bm{G}_{\tau_{1}})p(\bm{G}_{\tau_{1}}|\bm{G}_{\tau_{2}})\cdots p(\bm{G}_{\tau_{K-1}}|\bm{G}_{T}).$ (3)
@@
$\mathrm{d}\bm{G}_{t}^{(k)}=\mathbf{f}_{k,t}(\bm{G}_{t}^{(k)})\mathrm{d}t+g_{k,t}\mathrm{d}\bm{W}_{t},t\in[\tau_{k-1},\tau_{k}].\vskip-8.53581pt$ (4)
@@ 3.2 Guided Generation via Diffusion Bridge Process
Applying Doob’s $h$-transform (Doob and Doob, 1984) to the GOU process under the terminal condition $\bm{\mu}=\bm{G}_{\tau_{k}}$, we can derive the GOU bridge process as follows (detailed derivation of the bridge process provided in Sec. A.1):
@@
where $\omega(t)$ is a positive weighting function, and $c_{1},c_{2}$ control the relative importance of vertices and spectrum. The training procedure is detailed in Alg. 1 in App. C.
@@
To achieve this, we introduce a unified score network that explicitly integrates node and spectral representations. As illustrated in Fig. 7 in the Appendix, the network comprises two different graph processing modules: a standard graph convolution network (GCN) (Kipf and Welling, 2017) for local feature aggregation and a graph transformer network (ATTN) (Dwivedi and Bresson, 2021; Vignac et al., 2023) for global information extraction. The outputs of these modules are fused with time information through a Feature-wise Linear Modulation (FiLM) layer (Perez et al., 2018), and the resulting representations are concatenated to form a unified hidden embedding. This hidden embedding is further processed by separate multilayer perceptrons (MLPs) to produce predictions for $\nabla_{\bm{X}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$ and $\nabla_{\bm{\Lambda}}\log p(\bm{G}_{t}|\bm{G}_{\tau_{k}})$, respectively. It is worth noting that our score network is permutation equivalent, as each component of our model avoids any node ordering-dependent operations. Our model is detailed in Sec. C.3.